Assume that U is a plane. Find out whether or not the following vectors lie in U:

(10.1) ~u =< 3.8, 1 >, ~v =< −4, 1, 1 > and w~ = −~v

(10.2) ~u =< 3.8, 1 >, ~v =< −4, 1, 1 > and w~ = ~u − ~v


1
Expert's answer
2021-06-08T18:34:38-0400

1) The three vectors are coplanar if their scalar triple product is zero.


2) The three vectors are coplanar if they are linearly dependent.


(10.1)


"\vec u\cdot(\vec v\times \vec w)=\begin{vmatrix}
 3 & 8 & 1 \\
 -4 & 1 & 1 \\
 4 & -1 & -1
\end{vmatrix}"

"=3\begin{vmatrix}
 1 & 1 \\
 -1 & -1
\end{vmatrix}-8\begin{vmatrix}
 -4 & 1 \\
 4 & -1
\end{vmatrix}+1\begin{vmatrix}
 -4 &1 \\
 4 & -1
\end{vmatrix}"

"=3(0)-8(0)+0=0"

The following vectors lie in U.


 Since the vector "\vec u, \vec v" and "-\vec v" are linearly dependent, they lie in U.


(10.2)


"\vec u\cdot(\vec v\times \vec w)=\begin{vmatrix}
 3 & 8 & 1 \\
 -4 & 1 & 1 \\
 7 & 7 & 0
\end{vmatrix}"

"=3\begin{vmatrix}
 1 & 1 \\
 7 & 0
\end{vmatrix}-8\begin{vmatrix}
 -4 & 1 \\
 7 & 0
\end{vmatrix}+1\begin{vmatrix}
 -4 &1 \\
 7 & 7
\end{vmatrix}"

"=3(-7)-8(-7)+(-35)=0"

The following vectors lie in U.


 Since the vector "\vec u, \vec v" and "\vec u-\vec v" are linearly dependent, they lie in U.